6. fx) is a quadratic function, and go) is a linear fiunction for which 2 C-oo,-3) and (4,oo). Determine a possible equation set for fx) and gx). f (x) g(x) on the intervals of
Answer #1

Take f(x) x2 + bx + c and g(x) = x + d g(x) r+d Now this hold if P(x) = 2.2 + (b-1)x + c-d-1 has no zeroes. D=(1-1)2-4(c-d-1) < 0 Take d =-1